# On the Non-Orientable $3$- and $4$-Genera of a Knot: Connections and Comparisons

### v1

@@ 2.4. Essential Surfaces
We briefly recall some definitions of the required tools. Suppose $Y$ is a $3$-manifold with boundary and $F$ is a properly embedded surface in $Y$. Below, we will only consider when $Y$ is the exterior of a knot $K$, i.e. when $Y$ is $E(K)=S^{3}\setminus\nu(K)$. Choosing a meridian $\mu$ and Seifert-framed longitude $\lambda$ for $\partial\nu(K)$, we write $\partial F]=a[\mu]+b[\lambda]$. The boundary slope $s(F)$ is $\frac{a}{b}\in\mathbb{Q}\cup\{\frac{1}{0}\}$. This agrees with the definition of boundary slope for spanning surfaces, as for a spanning surface, $b=1$ and $a$ measures the linking number with respect to the Seifert framing. As before, the normal Euler number $e_{3}(F)$ is the negation of the boundary slope. We extend $\Gamma^{\pm}$ to any surface properly embedded in the exterior of a knot.
We follow [24] in defining a compressing disk for $F$ to be an embedded disk $D\subset Y$ with $D\cap F=\partial D$. The surface $F$ is incompressible if for each compressing disk $D$, there is a disk $D^{\prime}\subset F$ with $\partial D^{\prime}=\partial D$. Similarly, a $\partial$-compressing disk of $F$ is an embedded disk $D\subset Y$ with $D\cap F=\alpha$ and $D\cap\partial Y=\beta$, where $\alpha\cup\beta=\partial D$ and $\alpha\cap\beta=S^{0}$. The surface $F$ is $\partial$-incompressible if for each $\partial$-compressing disk $D$, there is a disk $D^{\prime}\subset F$ with $\partial D^{\prime}=\alpha\cup\beta^{\prime}$ with $\beta^{\prime}\subset\partial F$. A compression along a compressing or $\partial$-compressing disk is a surgery along the compressing or $\partial$- compressing disk. Finally, we say that a surface that is incompressible and $\partial$-incompressible, or is a sphere not bounding a ball, or is a disk that is not $\partial$-parallel, is essential.
@@
If $F$ is a properly embedded surface in $E(K)$, we set the notation $\tilde{W}_{F}=\tilde{W}_{(e_{3}(F),b_{1}(F))}$. Now let $\mathcal{F}(K)$ denote the set of all incompressible and $\partial$-incompressible surfaces in $E(K)$, which is a finite set by [22]. Define the incompressible wedge of $K$ to be
$\tilde{W}(K)=\bigcup_{F\in\mathcal{F}(K)}\tilde{W}_{F}.$
We may now state the main structural result in this section, which constrains the non-orientable $3$-dimensional geography of a knot $K$ using knowledge of the incompressible and $\partial$-incompressible surfaces in its exterior. This should be thought of as an analogue of Corollary 2.3.
@@
$\Gamma^{\pm}(F)\geq\min\{\Gamma^{\pm}(F^{\prime})\;:\;F^{\prime}\in\mathcal{F}(K)\}.$
@@ 3. Fillings in $4$ Dimensions
Our next step is to describe the framework necessary to investigate the non-orientable $4$-genus of a knot. In parallel to the $3$-dimensional setting, we first review the normal Euler number and Allen’s geography problem in Section 3.1. Next, in Section 3.2, we revisit pinch surfaces in $4$ dimensions. We finish the section by describing lower bounds on the non-orientable $4$-genus in Sections 3.3 and 3.4, including a proof of Theorem 1.2.
@@ 3.1. Normal Euler Number and $4$-Dimensional Geography
In the case where the cobordism $F$ arises from attaching a flat band to a diagram $D_{0}$ to obtain $D_{1}$, the normal Euler number may be computed using the blackboard framing in the formula in [38, Lemma 4.2]:
@@
That the three- and four-dimensional geographies of the trefoil are identical follows from our work from Section 2 and Proposition 3.1 in [3]. The $5_{2}$ knot, however, provides an example where the $3$-dimensional geography is a strict subset of the $4$-dimensional geography. In Figure 10, we display a filling $F$ of the $5_{2}$ knot arising from a flat band move. Using Equation (3.2), we compute that $e(F)=2$ and $b_{1}(F)=1$. Recalling Example 2.6, we see that the point $(2,1)\in R_{4}(K)$, but $(2,1)\not\in R_{3}(K)$.
@@
Figure 10. A Möbius strip filling of the $5_{2}$ knot with normal Euler number 2 arises from an isotopy, an attachment of a flat band, and finally an isotopy to the unknot.
3.2. Construction: Pinch Surfaces, Revisited
The pinch surface construction for torus knots described in Section 2.3 extends to finding fillings in $B^{4}$ — in fact, that was the design of the original description in [5]. For any positive, relatively prime $(p,q)$, use the pinch move to create a non-orientable cobordism in successive layers of $B^{4}$ from $T(p,q)$ down to the first instance of $T(r,1)$. Since $T(r,1)$ is an unknot, we can cap it off with a disk in $B^{4}$, yielding a filling surface $F_{4}(p,q)$.
Lemma 3.4 (Lemma 2.3 of [39]).
If $T(p,q)$ is transformed into $T(r,1)$ in the construction of $F_{4}(p,q)$, then
e(F(p,q))=rpq.e(F(p,q))=r-pq.
@@ 3.3. Signature and the Gordon-Litherland Bound
$\langle a,b\rangle_{F}=\operatorname{lk}(\alpha,p^{-1}\beta).$
This is a well-defined symmetric bilinear pairing [19], and we write $\sigma(F)$ for its signature. The signature of the pairing $\langle,\rangle_{F}$ is related to the signature of $K$ and the normal Euler number of $F$.
@@
$\sigma(K)=\sigma(F)+\frac{e(F)}{2}.$
@@
hence justifying taking the minimum in the definition of the Euler-normalized non-orientable genus.
The Gordon-Litherland inequality restricts the $3$- and $4$-dimensional geography of a knot $K$ to lie in the wedge $W_{\sigma}(K)$ based at $(2\sigma(K),0)$; see Figure 11. As noted in the introduction, the Euler-normalized non-orientable genera measure the minimum vertical distance from $R_{*}(K)$ to the boundary of $W_{\sigma}(K)$.
@@
Figure 11. The wedges $W_{\sigma}(K)$ and $W_{\upsilon}(K)$ restrict the $4$-dimensional (and hence 33-dimensional) non-orientable geography of a knot $K$. The bound in Corollary 3.8 is the height of the vertex of $W_{\sigma}(K)\cap W_{\upsilon}(K)$, while the vertical distances between the edges of the wedges yield the bounds in Theorem 1.2. The values for $\sigma$ and $\upsilon$ are taken from $T(4,3)$.
@@ 3.4. Further Bounds
The standard next step is to combine Theorems 3.6 and 3.7 to obtain a lower bound on the non-orientable $4$-genus.
@@
We are now ready to prove Theorem 1.2. The bound in Corollary 3.8 can be visualized as the height of the base of the wedge $W_{\sigma}(K)\cap W_{\upsilon}(K)$; see Figure 11 once again. Inspired by this visualization, we may also extract a lower bound on the Euler-normalized non-orientable $4$-genus, as the quantity $|2\upsilon(K)-\sigma(K)|$ measures the vertical distance between the boundaries of $W_{\sigma}(K)$ and $W_{\upsilon}(K)$. Theorem 1.2 follows.
4. Euler-Normalized Genus Agreements: Alternating Knots and the Turaev Genus
We return to dimension $3$ to examine the Euler-normalized non-orientable genera of alternating knots (Section 4.1), which leads to a connection to the Turaev genus (Section 4.2). We also explore some examples of torus knots in Section 4.3, and call attention to the relationship between the $3$- and $4$-dimensional Euler-normalized non-orientable genera for the knots examined in this section.
4.1. Alternating Knots
The goal of this section is to prove Theorem 1.3, i.e. that a knot is alternating if and only if both of its stable non-orientable $3$-genera vanish. We first prove the forward implication using a combinatorial approach, and then prove both implications using a topological approach based on Greene’s characterization of alternating knots. We note that the ideas in the combinatorial approach closely match those in [10], where the genera of non-orientable spanning surfaces for alternating knots were related to degrees of the Jones polynomial.
4.1.1. Combinatorial Proof of Theorem 1.3
Let $F_{A}$ (resp. $F_{B}$) be the all-$A$ (resp. all-$B$) state surface of a reduced alternating diagram $D$ of $K$. It suffices to show that $\Gamma^{+}(F_{A})=0=\Gamma^{-}(F_{B})$. In particular, we will show that
(4.1) $b_{1}(F_{A})=\frac{e(F_{A})}{2}-\sigma(K).$
A similar proof works for $F_{B}$.
First, we set notation. The number of crossings of a diagram $D$ is denoted $\operatorname{cr}$, with $\operatorname{cr}_{+}$ (resp. $\operatorname{cr}_{-}$) the number of positive (resp. negative) crossings. For a state $s$, let $|s|$ denote the number of state circles. For the standard checkerboard shading of a reduced alternating diagram $D$, the state circles of $s_{A}$ (resp. of $s_{B}$) correspond to the white (resp. black) regions.
There are three main inputs into the proof. The first is a computation of the normal Euler and Betti numbers of the state surfaces $F_{A}$ and $F_{B}$.
Lemma 4.1.
For a diagram $D$ of a knot $K$,
$e(F_{A})$ $=2\operatorname{cr}_{-}$ $e(F_{B})$ $=-2\operatorname{cr}_{+}$
$b_{1}(F_{A})$ $=1-|s_{A}|+\operatorname{cr}$ $b_{1}(F_{B})$ $=1-|s_{B}|+\operatorname{cr}$
The proof follows from Lemma 2.4 and the construction of the state surfaces. Alternatively, the computations of the normal Euler numbers appear as Lemma 5 in [15].
The second input is a fact about the combinatorics of an alternating diagram.
Lemma 4.2.
A diagram $D$ of an alternating knot $K$ satisfies
$2=-\operatorname{cr}+|s_{A}|+|s_{B}|.$
The proof follows from $\chi(S^{2})=2$, the fact that a reduced alternating diagram yields a cell decomposition of $S^{2}$, and the observation that an $A$-split separates adjacent white regions.
The last input computes the signature of an alternating knot directly from its diagram.
Lemma 4.3 (Theorem 2 of [43]).
Given an alternating knot $K$ with reduced alternating diagram $D$, we have
$\sigma(K)=\frac{1}{2}\left(-\operatorname{wr}(D)+|s_{A}|-|s_{B}|\right).$
We are now ready to prove the relation in Equation (4.1), and hence the forward direction of Theorem 1.3. We compute as follows:
$\frac{1}{2}e(F_{A})-b_{1}(F_{A})$ $=\operatorname{cr}_{-}-1+|s_{A}|-\operatorname{cr}$ (Lemma 4.1)
$=-\operatorname{cr}_{+}+\frac{1}{2}\left(\operatorname{cr}-|s_{A}|-|s_{B}|\right)+|s_{A}|$ (Lemma 4.2)
$=-\sigma(K)$ (Lemma 4.3)
4.1.2. Topological Proof of Theorem 1.3
An alternative approach to the proof of Theorem 1.3 uses Greene’s topological characterization of alternating links [21], adjusted slightly to our setting. We say that a spanning surface $F$ is positive (resp. negative) definite if the Gordon-Litherland form $\langle,\rangle_{F}$ is positive (resp. negative) definite.
Theorem 4.4 ([21]).
A knot $K\subset S^{3}$ is alternating if and only if $K$ bounds both a positive and a negative definite spanning surface.
First, suppose $K$ is an alternating knot. By Theorem 4.4, $K$ bounds a positive definite surface $F_{+}$. Combining the positive definiteness with Proposition 3.5, we see that
$b_{1}(F_{+})=\sigma(F_{+})=\sigma(K)-\frac{e(F_{+})}{2}.$
It follows immediately that $\hat{\gamma}^{+}_{3}(K)=0$. Similarly, since $K$ also bounds a negative definite surface $F_{-}$, we have $\hat{\gamma}^{-}_{3}(K)=0$.
Conversely, the calculations above show that if $\hat{\gamma}^{\pm}_{3}(K)=0$, then $K$ must bound both a positive and a negative definite surface. Hence, by Theorem 4.4, $K$ must be alternating.
4.2. Relation to the Turaev Genus
The Turaev genus $g_{T}$ of a knot diagram $D$ measures the difference between the all-$A$ and the all-$B$ state circles of the diagram. Specifically, we construct the Turaev surface $F_{T}$ of $D$ by taking a copy of the all-$A$ state circles $s_{A}$ just above the plane of the diagram $D$ and a copy of the all-$B$ state circles $s_{B}$ just below $D$. Over the edges of $D$, connect the state circles with a rectangular strip; near the crossings of $D$, connect them with a saddle as in Figure 12. This creates a cobordism between $s_{A}$ and $s_{B}$. Finally, cap off each of the state circles with a disk. The Turaev genus of the diagram $D$ is the genus of the Turaev surface $F_{T}$, which may be computed as
(4.2) $g_{T}(D)=\frac{1}{2}\left(\operatorname{cr}+2-|s_{A}|-|s_{B}|\right).$
The Turaev genus $g_{T}(K)$ of a knot $K$ is the minimum of the Turaev genera over all possible diagrams of $K$. See the original paper [12], the survey [8], or [31] for further discussion.
[…]
Figure 12. Connecting the state circles over the edges and near the crossings of a diagram $D$ to create a cobordism between $s_{A}$ and $s_{B}$.
We are now ready for the proof of Theorem 1.4.
Proof of Theorem 1.4.
It suffices to show that, for any diagram $D$ of a knot $K$, we have
$g_{T}(D)=\frac{1}{2}\left(\Gamma^{+}(F_{A})+\Gamma^{-}(F_{B})\right).$
This relation follows directly from the definition of $\Gamma^{\pm}$, Lemma 4.1, and a short computation. ∎
The bound in Theorem 1.4 is not always realized, as discussed in the next section. For an adequate knot, however, we conjecture that the bound is sharp; this was the statement of Conjecture 1.6. Theorem 1.3 shows that the conjecture holds for all alternating knots, and Theorem 1.4 implies that the conjecture holds for all Turaev genus $1$ knots. In order to prove the full conjecture, however, we would need to understand all possible spanning surfaces of adequate knots or, alternatively, all possible essential surfaces in the exterior of the knot. We will demonstrate how the latter type of argument would work in Section 6.
4.3. Torus Knots
@@
Theorem 1.3 implies that the Euler-normalized non-orientable $3$- and $4$-dimensional non-orientable genera agree for alternating knots. The torus knot $T(4,3)$ displays similar behavior, even though its ordinary $3$- and $4$-dimensional genera differ. Further, as with alternating knots, the bound in Theorem 1.4 for $T(4,3)$ is sharp, though non-zero.
@@
    $\hat{\gamma}^{+}_{3}(T(4,3))=0=\hat{\gamma}^{+}_{4}(T(4,3)$ while $\hat{\gamma}^{-}_{3}(T(4,3))=2=\hat{\gamma}^{-}_{4}(T(4,3)$, and
@@
The first part of the proposition follows from the computation that $\gamma_{3}(T(4,3))=2$ and $\gamma_{4}(T(4,3))=1$; see [29, Example 5.3] or the entry for $8_{19}=T(4,3)$ in [34]. These values are realized using pinch surfaces, which have
@@
The proof of the second part starts with the computations $\sigma(T(4,3))=-6$ and $\upsilon(T(4,3))=-2$ using techniques from [30]. Thus, the bounds in Theorem 1.2 are sharp at the pinch surface $F_{4}(4,3)$, and hence $\hat{\gamma}^{+}_{4}(T(4,3))=0$ and $\hat{\gamma}^{-}_{4}(T(4,3))=2$. The computations for the second part of the proof are summarized in terms of geography in Figure 13.
Turning to spanning surfaces of $T(4,3)$ in $S^{3}$, on the positive side, the minimal Seifert surface $S$ has $e(S)=0$ and $b_{1}(S)=6$, and hence $\Gamma_{3}^{+}(\operatorname{tw}_{+}(S))=0$. It follows that $\hat{\gamma}^{+}_{3}(T(4,3))=0$. On the negative side, the pinch surface $F_{3}(4,3)$ has first Betti number $2$ and normal Euler number 12-12, and hence has $\Gamma^{-}_{3}(F_{3}(4,3))=2$. As the lower bounds in Theorem 1.2 also apply to $\hat{\gamma}^{\pm}_{3}$, it follows that $\hat{\gamma}^{-}_{3}(T(4,3))=2$ as well. Once again, see Figure 13.
@@
Figure 13. The 44-dimensional geography of the torus knot $T(4,3)$ is precisely $W_{F_{4}(p,q)}$, while the $3$-dimensional geography contains both $W_{F_{3}(p,q)}$ and $W_{\operatorname{tw}_{+}(S)}$ (and may contain more).
@@
The torus knot $T(4,3)$ is the first non-trivial example in the family of torus knots T(2k,2k1)T(2k,2k-1). Batson [5] showed that this family has unbounded $\gamma_{4}$, with $\gamma_{4}(T(2k,2k-1))=k-1$. Further, all elements in the family have $R_{4}=W_{\sigma}\cap W_{\upsilon}$ [39], which yields
$\hat{\gamma}^{-}_{4}(T(2k,2k-1))=0\quad\text{and}\quad\hat{\gamma}^{+}_{4}(T(2k,2k-1))=2k-2.$
Thus, we obtain the result in Corollary 1.5:
$g_{T}(T(2k,2k-1))$ $\geq\overline{\gamma}_{3}(T(2k,2k-1))$
$\geq\overline{\gamma}_{4}(T(2k,2k-1))$
$=k-1.$
We note that for T(2k,2k1)T(2k,2k-1), it is not clear if $\hat{\gamma}^{+}_{3}=\hat{\gamma}^{+}_{4}$; in fact, this seems unlikely to be the case, as the Seifert surface lies inside the wedge of the $3$-dimensional pinch surface for $k>1$.
5. Ordinary Genus Gaps: Pretzel Knots
At the end of the previous section, we saw that the torus knot $T(4,3)$ has a gap of $1$ between its non-orientable $3$- and $4$-genera. In fact, Jabuka and Van Cott [29] prove that the gap can be arbitrarily large on torus knots:
@@ 5. Ordinary Genus Gaps: Pretzel Knots
In this section, we provide a new family of examples for which the gap becomes arbitrarily large. These examples, which are a family of pretzel knots, have several interesting features. First, the new examples all bound Möbius bands in $B^{4}$ even though the orientable $4$-genus becomes arbitrarily large. Second, the gap between the Euler-normalized non-orientable $3$- and $4$-genera is bounded above by $2$, demonstrating that the Euler-normalized non-orientable gaps are independent of ordinary non-orientable gaps.

### v2

@@ 2.2. Essential Surfaces
We briefly recall some definitions of the required tools. Suppose $Y$ is a $3$-manifold with boundary and $F$ is a properly embedded surface in $Y$. Below, we will only consider when $Y$ is the exterior of a knot $K$, i.e. when $Y$ is $E(K)=S^{3}\setminus\nu(K)$. Choosing a meridian $\mu$ and Seifert-framed longitude $\lambda$ for $\partial\nu(K)$, we write $\partial F]=a[\mu]+b[\lambda]$. The boundary slope $s(F)$ is $\frac{a}{b}\in\mathbb{Q}\cup\{\frac{1}{0}\}$. This agrees with the definition of boundary slope for spanning surfaces, as for a spanning surface, $b=1$ and $a$ measures the linking number with respect to the Seifert framing. As before, the normal Euler number $e_{3}(F)$ is the negation of the boundary slope. We extend the Euler-normalized first Betti numbers $\Gamma^{\pm}$ to any surface properly embedded in the exterior of a knot.
We follow [20] in defining a compressing disk for $F$ to be an embedded disk $D\subset Y$ with $D\cap F=\partial D$. The surface $F$ is incompressible if for each compressing disk $D$, there is a disk $D^{\prime}\subset F$ with $\partial D^{\prime}=\partial D$. Similarly, a $\partial$-compressing disk of $F$ is an embedded disk $D\subset Y$ with $D\cap F=\alpha$ and $D\cap\partial Y=\beta$, where $\alpha\cup\beta=\partial D$ and $\alpha\cap\beta=S^{0}$. The surface $F$ is $\partial$-incompressible if for each $\partial$-compressing disk $D$, there is a disk $D^{\prime}\subset F$ with $\partial D^{\prime}=\alpha\cup\beta^{\prime}$ with $\beta^{\prime}\subset\partial F$. A compression along a compressing or $\partial$-compressing disk is a surgery on $F$ along the compressing or $\partial$- compressing disk. Finally, we say that a surface that is incompressible and $\partial$-incompressible, or is a sphere not bounding a ball, or is a disk that is not $\partial$-parallel, is essential.
@@
If $F$ is a properly embedded surface in $E(K)$, we set the notation $\tilde{W}_{F}=\tilde{W}_{(e_{3}(F),b_{1}(F))}$. Now let $\mathcal{F}_{\text{ess}}(K)$ denote the set of all incompressible and $\partial$-incompressible surfaces in $E(K)$, which is a finite set by [18]. Define the incompressible wedge of $K$ to be
$\tilde{W}(K)=\bigcup_{F\in\mathcal{F}_{\text{ess}}(K)}\tilde{W}_{F}.$
We may now state the main structural result in this section, which constrains the non-orientable $3$-dimensional geography of a knot $K$ using knowledge of the incompressible and $\partial$-incompressible surfaces in its exterior. This should be thought of as an analogue of Equation (2.2).
@@
$\Gamma^{\pm}(F)\geq\min\{\Gamma^{\pm}(F^{\prime})\;:\;F^{\prime}\in\mathcal{F}_{\text{ess}}(K)\}.$
@@ 3. Fillings in $4$ Dimensions
Our next step is to describe the framework necessary to investigate the non-orientable $4$-genus of a knot. In parallel to the $3$-dimensional setting, we first review the normal Euler number and Allen’s geography problem in Section 3.1. We finish the section by describing lower bounds on the non-orientable $4$-genus in Sections 3.2 and 3.3, including a proof of Proposition 1.5.
@@ 3.1. Normal Euler Number and $4$-Dimensional Geography
In the case where the cobordism $F$ arises from attaching a flat band to a diagram $D_{0}$ to obtain $D_{1}$, the normal Euler number may be computed using the blackboard framing in the formula in [37, Lemma 4.2]:
@@
That the three- and four-dimensional geographies of the trefoil are identical follows from our work from Section 2 and Proposition 3.1 in [3]. The $5_{2}$ knot, however, provides an example where the $3$-dimensional geography is a strict subset of the $4$-dimensional geography. In Figure 6, we display a filling $F$ of the $5_{2}$ knot arising from a flat band move. Using Equation (3.2), we compute that $e(F)=2$ and $b_{1}(F)=1$. Thus, we see that the point $(2,1)\in R_{4}(K)$. On the other hand, using the algorithm of Adams and Kindred [2], the point $(2,1)$ is not in $R_{3}(K)$.
@@
Figure 6. A Möbius strip filling of the $5_{2}$ knot with normal Euler number 2 arises from an isotopy, an attachment of a flat band, and finally an isotopy to the standard unknot.
@@ 3.2. Signature and the Gordon-Litherland Bound
$\langle a,b\rangle_{F}=\operatorname{lk}(\alpha,p^{-1}\beta),$
which yields a well-defined symmetric bilinear pairing [15]. The signature $\sigma_{gl}$ of the pairing $\langle,\rangle_{F}$ is related to the signature of $K$ and the normal Euler number of $F$.
@@
$\sigma(K)=\sigma_{gl}(F)+\frac{e(F)}{2}.$
In particular, as noted in the introduction, we can rewrite the normalization in $\Gamma^{\pm}(F)$ as $\sigma_{gl}(F)$ rather than $\sigma(K)-\frac{e(F)}{2}$.
@@
hence justifying taking the minimum in the definition of the $\pm$-spanning surface defect.
The Gordon-Litherland inequality restricts the $3$- and $4$-dimensional geography of a knot $K$ to lie in the wedge $W_{\sigma}(K)$ based at $(2\sigma(K),0)$; see Figure 7. As noted in the introduction, the spanning surface defects measure the minimum vertical distance from $R_{*}(K)$ to the boundary of $W_{\sigma}(K)$.
@@
Figure 7. The wedges $W_{\sigma}(K)$ and $W_{\upsilon}(K)$ restrict the $4$-dimensional (and hence 33-dimensional) non-orientable geography of a knot $K$. The bound in Corollary 3.7 is the height of the vertex of $W_{\sigma}(K)\cap W_{\upsilon}(K)$, while the vertical distances between the edges of the wedges yield the bounds in Proposition 1.5. The values for $\sigma$ and $\upsilon$ are taken from $T(4,3)$.
@@ 3.3. Further Bounds
The standard next step is to combine Theorems 3.5 and 3.6 to obtain a lower bound on the non-orientable $4$-genus.
@@
We are now ready to prove Proposition 1.5. The bound in Corollary 3.7 can be visualized as the height of the base of the wedge $W_{\sigma}(K)\cap W_{\upsilon}(K)$; see Figure 7 once again. Inspired by this visualization, we may also extract a lower bound on the Euler-normalized non-orientable $4$-genus, as the quantity $|2\upsilon(K)-\sigma(K)|$ measures the vertical distance between the boundaries of $W_{\sigma}(K)$ and $W_{\upsilon}(K)$. The proposition follows.
4. Motivation from Torus Knots
@@
Before tackling the main results of the paper, we take a motivational interlude to briefly discuss torus knots. Theorem 1.2 implies that $3$- and $4$-dimensional spanning surface defects agree for alternating knots. The torus knot $T(4,3)$ displays similar behavior, even though its ordinary $3$- and $4$-dimensional genera differ. Further, as with alternating knots, the bound in Theorem 1.3 for $T(4,3)$ is sharp, though non-zero.
@@
    $d^{+}_{3}(T(4,3))=0=d^{+}_{4}(T(4,3)$ while $d^{-}_{3}(T(4,3))=2=d^{-}_{4}(T(4,3)$, and
@@
The first part of the proposition follows from the computation that $\gamma_{3}(T(4,3))=2$ and $\gamma_{4}(T(4,3))=1$; see [28, Example 5.3] or the entry for $8_{19}=T(4,3)$ in [32]. These values are realized using the $3$- and $4$-dimensional pinch surfaces $F_{*}(4,3)$ defined in [28] and [5], respectively. Using [38, Lemma 2.3], we may compute that
@@
The proof of the second part starts with the computations $\sigma(T(4,3))=-6$ and $\upsilon(T(4,3))=-2$ using techniques from [29]. Thus, the bounds in Proposition 1.5 are sharp at the pinch surface $F_{4}(4,3)$, and hence $d^{+}_{4}(T(4,3))=0$ and $d^{-}_{4}(T(4,3))=2$. The computations for the second part of the proof are summarized in terms of geography in Figure 8.
Turning to spanning surfaces of $T(4,3)$ in $S^{3}$, on the positive side, the minimal Seifert surface $S$ has $e(S)=0$ and $b_{1}(S)=6$, and hence $\Gamma_{3}^{+}(\operatorname{tw}_{+}(S))=0$. It follows that $d^{+}_{3}(T(4,3))=0$. On the negative side, the pinch surface $F_{3}(4,3)$ has first Betti number $2$ and normal Euler number 12-12, and hence has $\Gamma^{-}_{3}(F_{3}(4,3))=2$. As the lower bounds in Proposition 1.5 also apply to $d^{\pm}_{3}$, it follows that $d^{-}_{3}(T(4,3))=2$ as well. Once again, see Figure 8.
@@
Figure 8. The $4$-dimensional geography of the torus knot $T(4,3)$ is precisely $W_{F_{4}(p,q)}$, while the $3$-dimensional geography contains both $W_{F_{3}(p,q)}$ and $W_{\operatorname{tw}_{+}(S)}$ (and may contain more).
@@
The proof of the first part of the proposition shows that $T(4,3)$ has a gap of $1$ between its non-orientable $3$- and $4$-genera. In fact, Jabuka and Van Cott [28] prove that the gap can be arbitrarily large on torus knots:
@@
The torus knot $T(4,3)$ is also the first non-trivial example in the family of torus knots T(2k,2k1)T(2k,2k-1). Batson [5] showed that this family has unbounded $\gamma_{4}$, with $\gamma_{4}(T(2k,2k-1))=k-1$. Further, all elements in the family have $R_{4}=W_{\sigma}\cap W_{\upsilon}$ [38], which yields
$d^{-}_{4}(T(2k,2k-1))=0\quad\text{and}\quad d^{+}_{4}(T(2k,2k-1))=2k-2.$
We note that for T(2k,2k1)T(2k,2k-1), it is not clear if $d^{+}_{3}=d^{+}_{4}$; in fact, this seems unlikely to be the case, as the Seifert surface lies inside the wedge of the $3$-dimensional pinch surface for $k>1$.
5. Ordinary Genus Gaps for Pretzel Knots
In this section, we provide a new family of examples for which the gap between the non-orientable $3$- and $4$-genera becomes arbitrarily large. These examples, which are a family of pretzel knots, have several interesting features. First, the new examples all bound Möbius bands in $B^{4}$ even though the orientable $4$-genus becomes arbitrarily large. Second, the gap between the Euler-normalized non-orientable $3$- and $4$-genera is bounded above by $2$, demonstrating that the spanning surface defect gaps are independent of ordinary non-orientable gaps.